# 00footnotetext: Shared first author; Shared last author

### v1

@@ Appendix A Contributions
Alon Albalak, Hritik Bansal, Yair Carmon, Achal Dave, Alex Dimakis, Alex Fang, Samir Gadre, Etash Guha, Reinhard Heckel, Maor Ivgi, Sedrick Keh, Jeffrey Li, Niklas Muennighoff, Sarah Pratt, Ludwig Schmidt, Vaishaal Shankar, Georgios Smyrnis
@@ Appendix C Benchmark rules
The defining characteristic of entries in the filtering track is that they form the dataset by applying a processing pipeline on the subset of DCLM-Pool corresponding to the chosen compute scale (see Table˜1) without including any external data. The rationale behind this requirement is twofold. First, the size and quality of initial data for filtering affects both the processing cost and the quality of the processed dataset. By fixing the initial dataset we level the playing field and allow comparison to focus on core curation techniques. Second, we wish to encourage the development of methods potentially relevant even at frontier-model scale. Using the 7B-2x pool (containing roughly 16T tokens) for the 400M-1x compute scale (requiring roughly 8B tokens for training) would allow filtering strategies that keep less than 0.1% of the data and cannot scale to generating a trillion-token dataset.
@@ Appendix D Tooling
For the construction of our pool, we download WARC files from Common Crawl, and process them via resiliparse, we do this by streaming data directly from S3 to EC2 using the Ray data processing framework. This is the starting point for our data processing pipeline. For the dataset released to participants, we release various sizes of DCLM-Pool, that we make available for download. For details on the data, see Appendix˜E.
@@
We base our training code on OpenLM [68], and provide configuration files for each of our scales. We also provide scripts that train models using each configuration, and produce json files that describe a trained model in detail. For further training details, see Appendix˜F.
@@
All of our results, including data processing, model training, evaluations, and plots included in this paper, are reproducible using our open-source framework and the recipes in https://datacomp.ai/dclm. We list compute requirements for our code in Appendix˜Q.
@@ Appendix E DCLM-Pool
DCLM-Pool was collected by taking all 5.1M Common Crawl WARC dumps from 2008 to 2022 (inclusive) and extracting text from the html using the resiliparse framework. We opted to omit 2023 and above to prevent large amounts of language model generated text from polluting our datasets and to provide a hold out for future use. We release DCLM-Pool on HuggingFace with CC-BY-4 license. We release DCLM-Pool as a set of .jsonl files similar to Dolma-V1 and RedPajama. We provide the fields that are in the .jsonl in Table˜10. The entire pool is 5.1M gzip compressed .jsonl\ files, and 340TB compressed on disk. The use of this dataset is also subject to CommonCrawl’s Terms of Use: https://commoncrawl.org/terms-of-use.
@@
We note that Common Crawl includes raw data as collected from the web without filtering. While some of our pools, such as DCLM-baseline, underwent some filtering of malicious URLs, none have had any special treatment for PII and sensitive content to preserve representativeness of the raw data. For a more complete discussion on PII and consent regarding our pools, see Appendix˜S.
@@ Appendix F Training details
We detail the hyperparameters for our models in Table˜11. For the 400M-1x and 1B-1x, we follow hyperparameters from [56], which were tuned to optimize perplexity on a validation set containing tokens from recent arXiv papers, the OpenLM codebase itself, and news articles. For the 1B-1x scale, we also investigated alternative hyperparameters in Table˜13, and find the hyperparameters from [56] perform best. For the 7B-1x and 7B-2x, we used a higher learning rate, and a lower weight decay, guided by the hyperparameter sweep in Table˜12. We use a cooldown of 3ee-5 for all experiments. For Table˜2, we trained with a lower learning rate following [56] as these experiments were performed before our sweep. Specifically, we used a learning rate of 3ee-4 and weight decay of 0.33.
@@
400M-1x 24 8 1,024 128 2,000 3ee-3 0.033 1ee-4 512
1B-1x 24 16 2,048 128 5,000 3ee-3 0.033 1ee-4 256
7B-1x, 7B-2x 32 32 4,096 128 5,000 2ee-3 0.05 5e-6 2,048
Table 12: Learning rate and weight decay sweep (7B-1x scale). We evaluated the impact of learning rate and weight decay on an earlier iteration of DCLM-baseline. Based on this sweep, we specify the settings for Table˜11 for the 7B-1x and 7B-2x scales.
@@ Appendix G Evaluation details
Given the multiple ways of evaluating accuracy [51], we conducted a miniature study using the LightEval evaluation framework [54]. Notably, under this framework, we are able to achieve scores above random (25%) for 0-shot MMLU for 1B models by considering the log-probabilities of entire answer passages as opposed to single letters. The 1B Hugging Face model trained on FineWeb-Edu [97] has shown to work well on this, so we wanted to more closely examine how LightEval evaluation scores correlate with evaluation scores from LLM Foundry. We present our findings in Figure˜5.
@@
The key difference between LightEval and LLM Foundry for multiple choice tasks like MMLU is that LightEval considers the log probabilities of entire answer sequences, whereas LLM Foundry only considers log probabilities of single letters. Nonetheless, Figure˜5 shows a positive correlation between the two evaluation frameworks on MMLU 0-shot accuracy.
In Figure˜5 we were able to reproduce the MMLU scores reported in the FineWeb-Edu blog [97]. Notably, we found that LightEval indeed gave MMLU scores above random for the 1B scales, whereas in LLM Foundry, all the 1B models have accuracies around 0.25. At larger scales, however, the LightEval scores for the models become quite cramped together, which may make it more difficult to compare models and may make the comparisons more susceptible to noise. For example, the models Gemma-7B, Llama3-8B, and Mistral-7B all have scores between 0.43 and 0.44 in LightEval, while their scores range from 0.56 to 0.62 for LLM Foundry. We also see that FineWeb-Edu 7B-2x and DCLM 7B-2x perform quite similarly in LightEval, but DCLM-7B is better by close to 10 points in LLM Foundry. In conclusion, we believe that LightEval can be potentially a good choice when evaluating smaller models, but other frameworks like LLM Foundry could give clearer signals when comparing larger models.
@@ Appendix H Hyperparameter study
A potential concern is that the training recipe can change conclusions about which dataset is optimal for training, due to interaction between training hyperparameters and dataset distributions. To address this confounder, we show in Table˜13 that orderings between datasets are preserved for various combinations of weight decay and learning rate. Moreover we find that performance gains from optimal hyper-parameter choice and dataset design tend to be orthogonal and complement each other. We illustrate this effect in Table˜14.
@@ Appendix I Model-based quality filters
We presented results for several different model-based quality filters in Section˜4.4. In this section, we describe their implementations in further detail, focusing especially on fastText classifiers which were our method of choice for DCLM-baseline.
@@
We use the supervised fastText package from Joulin et al. [79] to train models to classify between chosen “high-quality” reference data which are given positive labels, and web-crawled data which are given negative labels. We then apply these classifiers to score each document from the pool we wish to filter, taking the predicted probability of the positive label as the score and computing a percentile-based threshold. In terms of training hyperparameters, we mostly used the default choices from the fastText package; the only hyperparameter change that we tried was to expand the feature space from unigrams only to both unigrams and bigrams (via setting the wordNgrams argument to be 2 instead of the default 1). This helped improve the quality of downstream filtered datasets, as shown in Table˜15 (which extends Table˜5 from Section˜4.4).
@@
The bulk of our experimentation for training fastText models focused on constructing their underlying training sets, specifically the positively labeled reference data. For each experiment, we fixed the size of the training set to be 400K examples (i.e., 200K positive, 200K negative). The negatively labeled examples were sampled randomly from a set of documents that came from an earlier (smaller) version of our RefinedWeb reproduction. This version used trafilatura as the extractor instead of resiliparse, which we hypothesize might actually help for training the filtering model; as shown in Appendix˜J, trafilatura more aggressively removes boilerplate content that may appear in many pages (especiall from the same website). This type of content, if left in, may lead to the fastText models over-relying on these “spurious” features instead of the main contents of the page. For the positively labeled reference data, we tried several different sources, some of which involved further pre-processing:
@@
We also examined other quality filters, though found none as effective as the fastText methods described above, as shown in Table˜4. We now provide further details for some of these baselines.
@@
An intuitively promising, but ultimately unfruitful approach was to consider page centrality metrics such as PageRank and Harmonic centrality metrics, with the idea that more "central" web text would yield higher quality data. We collected PageRank metrics from Common Crawl’s host level webgraph dataset [^3] https://commoncrawl.org/web-graphs and omitted any hosts that did not appear in the crawl. Next we partitioned our RefinedWeb reproduction into quintiles based on their PageRank score and trained several models at the 1B-1x scale. These results are collated in Table˜16, but unfortunately no quintile performed better than a pool sampled from the union of all quintiles.
@@
        Evaluate the following paragraph for LLM pre-training suitability:        - Is it well-structured or contains useful examples to natural texts?        - Does it offer insights, useful facts or relevant information?        {internallinenumbers*} - Does it teach how to comply with open-ended tasks such as writing letters, poems, emails etc.?        - Is it free from harmful content?              {internallinenumbers*} If most criteria are met based on the content below, indicate ’yes’ for suitable. Otherwise, indicate ’no’ for unsuitable.              ### <input> ###              Is it suitable for LLM pre-training? OPTIONS:        - yes        - no
where <input> is replaced with the document tokens, clipped if too long, and appended with ‘‘…[The rest of the paragraph is omitted]’’ in such cases. While this method worked slightly better than random sampling from our pool (see Table˜4), it significantly underperformed compared to our fastText experiments. Considering the high costs associated with applying it at scale, we did not perform this experiment on a larger scale.
@@
Following the success of the different deduplication methods we used (Appendix˜K), we studied the effect of Semantic deduplication as proposed by Abbas et al. [1]. In this approach, the authors propose embedding the documents using pre-trained language models, clustering them using k-means, and removing all but one document from each group of closely related documents to encourage diversity in the dataset. We began by embedding each document in a pool of approximately 100 million documents (following Abbas et al. [1]’s best practices) with BGE-base [172]. We then used faiss-GPU [77] to perform spherical k-means clustering, with 20 iterations and $K=11000$. We sampled documents after discarding 25% of the data. As seen in Table˜4, this intervention only negatively impacted the trained model. We hypothesize that the model used for embedding has a significant impact on the outcomes of this method. However, due to the large computational overhead when scaled, making it infeasible, we opted to rely on the deduplication methods outlined in Appendix˜K and leave this line of research for future work.
@@
We compute basic summary statistics for each extractor based on a sample of 10 WARC files (corresponding to 900K individual pages), presenting the results in Table˜17. Notably, both resiliparse and trafilatura result at least 2x shorter documents on average compared to WET files. As shown in the examples in Section˜J.2, WET files indeed contain many additional lines with seemingly little value for pre-training (e.g. navigation bars, boilerplate notices, copyright statements). trafilatura and resiliparse trim most of these lines out, with the former being more strict about doing so. Between the two, resiliparse still keeps in about 10% more text; some of this additional text may provide useful content such as section titles and dates for articles. In terms of runtime, the two are much farther apart, with resiliparse being roughly 8x faster.
@@
        [Trafilatura]              {internallinenumbers*} HERE is a sampling of some of the better antiques and flea markets around the United States.        Two or Three Times a Year        {internallinenumbers*} BRIMFIELD Route 20, Brimfield, Mass. 01010; 413-245-3436. Second weekend of May and July, and the second weekend after Labor Day.        {internallinenumbers*} RENNINGER’S OUTDOOR EXTRAVAGANZA Noble Street, Kutztown, Pa.; 717-385-0104. Thursday, Friday and Saturday of the last weekend of April, June, September.        {internallinenumbers*} FARMINGTON ANTIQUES WEEKEND Farmington Polo Grounds, Town Farm Road, Farmington, Conn. 06032; 508-839-9735. Starting Wednesday before shows open; 203-677-7862. June 9-10 and Sept. 1-2.        Monthly        {internallinenumbers*} ANN ARBOR ANTIQUES MARKET, P.O. Box 1512, Ann Arbor, Mich. 48106; 313-662-9453. May through October, third Sunday.        {internallinenumbers*} Continue reading the main storyKANE COUNTY FLEA MARKET, Kane County Fairgrounds, P.O. Box 549, St. Charles, Ill. 60174; 708-377-2252. Year-round, first weekend.        {internallinenumbers*} THE METROLINA EXPO, 7100 Statesville Road, Charlotte, N.C. 28213; 704-596-4643. Year-round, first weekend of every month.        {internallinenumbers*} SPRINGFIELD ANTIQUE SHOW AND FLEA MARKET, Clark County Fairgrounds, Route 41, Springfield, Ohio, 45501; 513-325-0053. Year-round, third weekend.        Weekly        {internallinenumbers*} BAKERSFIELD SWAP-O-RAMA, 4501 Wible Road, Bakersfield, Calif. 93313; 805-831-9342. Saturday and Sunday.        {internallinenumbers*} LAMBERTVILLE ANTIQUE MARKET, Route 29, Lambertville, N.J. 08530. Weekend number: 609-397-0456. Weekday: 215-752-4485, between 5 and 7 P.M. Market on Saturday and Sunday.        {internallinenumbers*} ATLANTA FLEA MARKET AND ANTIQUE CENTER, 5360 Peachtree Industrial Boulevard, Chamblee, Ga. 30341; 404-458-0456. Friday, Saturday and Sunday.        Continue reading the main story
@@
        [Resiliparse]              {internallinenumbers*} This is a digitized version of an article from The Times’s print archive, before the start of online publication in 1996. To preserve these articles as they originally appeared, The Times does not alter, edit or update them.              {internallinenumbers*} Occasionally the digitization process introduces transcription errors or other problems. Please send reports of such problems to archive_feedback@nytimes.com.              May 10, 1990, Page 00006 The New York Times Archives              {internallinenumbers*} HERE is a sampling of some of the better antiques and flea markets around the United States.              Two or Three Times a Year              {internallinenumbers*} BRIMFIELD Route 20, Brimfield, Mass. 01010; 413-245-3436. Second weekend of May and July, and the second weekend after Labor Day.              {internallinenumbers*} RENNINGER’S OUTDOOR EXTRAVAGANZA Noble Street, Kutztown, Pa.; 717-385-0104. Thursday, Friday and Saturday of the last weekend of April, June, September.              {internallinenumbers*} FARMINGTON ANTIQUES WEEKEND Farmington Polo Grounds, Town Farm Road, Farmington, Conn. 06032; 508-839-9735. Starting Wednesday before shows open; 203-677-7862. June 9-10 and Sept. 1-2.              Monthly              {internallinenumbers*} ANN ARBOR ANTIQUES MARKET, P.O. Box 1512, Ann Arbor, Mich. 48106; 313-662-9453. May through October, third Sunday.              Continue reading the main story              {internallinenumbers*} KANE COUNTY FLEA MARKET, Kane County Fairgrounds, P.O. Box 549, St. Charles, Ill. 60174; 708-377-2252. Year-round, first weekend.              {internallinenumbers*} THE METROLINA EXPO, 7100 Statesville Road, Charlotte, N.C. 28213; 704-596-4643. Year-round, first weekend of every month.              {internallinenumbers*} SPRINGFIELD ANTIQUE SHOW AND FLEA MARKET, Clark County Fairgrounds, Route 41, Springfield, Ohio, 45501; 513-325-0053. Year-round, third weekend.              Weekly              {internallinenumbers*} BAKERSFIELD SWAP-O-RAMA, 4501 Wible Road, Bakersfield, Calif. 93313; 805-831-9342. Saturday and Sunday.              {internallinenumbers*} LAMBERTVILLE ANTIQUE MARKET, Route 29, Lambertville, N.J. 08530. Weekend number: 609-397-0456. Weekday: 215-752-4485, between 5 and 7 P.M. Market on Saturday and Sunday.              {internallinenumbers*} ATLANTA FLEA MARKET AND ANTIQUE CENTER, 5360 Peachtree Industrial Boulevard, Chamblee, Ga. 30341; 404-458-0456. Friday, Saturday and Sunday.              Continue reading the main story

### v2

@@ Appendix A Contributions
Alon Albalak, Hritik Bansal, Yair Carmon, Achal Dave, Alexandros G. Dimakis, Alex Fang, Samir Gadre, Etash Guha, Reinhard Heckel, Maor Ivgi, Sedrick Keh, Jeffrey Li, Niklas Muennighoff, Sarah Pratt, Ludwig Schmidt, Vaishaal Shankar, Georgios Smyrnis
@@ Appendix C Benchmark rules
The defining characteristic of entries in the filtering track is that they form the dataset by applying a processing pipeline on the subset of DCLM-Pool corresponding to the chosen compute scale (see Table 1) without including any external data. The rationale behind this requirement is twofold. First, the size and quality of initial data for filtering affects both the processing cost and the quality of the processed dataset. By fixing the initial dataset we level the playing field and allow comparison to focus on core curation techniques. Second, we wish to encourage the development of methods potentially relevant even at frontier-model scale. Using the 7B-2x pool (containing roughly 16T tokens) for the 400M-1x compute scale (requiring roughly 8B tokens for training) would allow filtering strategies that keep less than 0.1% of the data and cannot scale to generating a trillion-token dataset.
@@ Appendix D Tooling
For the construction of our pool, we download WARC files from Common Crawl, and process them via resiliparse, we do this by streaming data directly from S3 to EC2 using the Ray data processing framework. This is the starting point for our data processing pipeline. For the dataset released to participants, we release various sizes of DCLM-Pool, that we make available for download. For details on the data, see Appendix E.
@@
We base our training code on OpenLM [70], and provide configuration files for each of our scales. We also provide scripts that train models using each configuration, and produce json files that describe a trained model in detail. For further training details, see Appendix F.
@@
All of our results, including data processing, model training, evaluations, and plots included in this paper, are reproducible using our open-source framework and the recipes in https://datacomp.ai/dclm. We list compute requirements for our code in Appendix Q.
@@ Appendix E DCLM-Pool
DCLM-Pool was collected by taking all 5.1M Common Crawl WARC dumps from 2008 to 2022 (inclusive) and extracting text from the html using the resiliparse framework. We opted to omit 2023 and above to prevent large amounts of language model generated text from polluting our datasets and to provide a hold out for future use. We release DCLM-Pool on HuggingFace with CC-BY-4 license. We release DCLM-Pool as a set of .jsonl files similar to Dolma-V1 and RedPajama. We provide the fields that are in the .jsonl in Table 10. The entire pool is 5.1M gzip compressed .jsonl\ files, and 340TB compressed on disk. The use of this dataset is also subject to CommonCrawl’s Terms of Use: https://commoncrawl.org/terms-of-use.
@@
We note that Common Crawl includes raw data as collected from the web without filtering. While some of our pools, such as DCLM-baseline, underwent some filtering of malicious URLs, none have had any special treatment for PII and sensitive content to preserve representativeness of the raw data. For a more complete discussion on PII and consent regarding our pools, see Appendix S.
@@ Appendix F Training details
[…]
@@
[…]
Table 12: Learning rate and weight decay sweep (7B-1x scale). We evaluated the impact of learning rate and weight decay on an earlier iteration of DCLM-baseline. Based on this sweep, we specify the settings for Table 11 for the 7B-1x and 7B-2x scales.
@@ Appendix G Evaluation details
Given the multiple ways of evaluating accuracy [53], we conducted a miniature study using the LightEval evaluation framework [56]. Notably, under this framework, we are able to achieve scores above random (25%) for 0-shot MMLU for 1B models by considering the log-probabilities of entire answer passages as opposed to single letters. The 1B Hugging Face model trained on FineWeb-Edu [100] has shown to work well on this, so we wanted to more closely examine how LightEval evaluation scores correlate with evaluation scores from LLM Foundry. We present our findings in Figure 5.
@@
The key difference between LightEval and LLM Foundry for multiple choice tasks like MMLU is that LightEval considers the log probabilities of entire answer sequences, whereas LLM Foundry only considers log probabilities of single letters. Nonetheless, Figure 5 shows a positive correlation between the two evaluation frameworks on MMLU 0-shot accuracy.
In Figure 5 we were able to reproduce the MMLU scores reported in the FineWeb-Edu blog [100]. Notably, we found that LightEval indeed gave MMLU scores above random for the 1B scales, whereas in LLM Foundry, all the 1B models have accuracies around 0.25. At larger scales, however, the LightEval scores for the models become quite cramped together, which may make it more difficult to compare models and may make the comparisons more susceptible to noise. For example, the models Gemma-7B, Llama3-8B, and Mistral-7B all have scores between 0.43 and 0.44 in LightEval, while their scores range from 0.56 to 0.62 for LLM Foundry. We also see that FineWeb-Edu 7B-2x and DCLM 7B-2x perform quite similarly in LightEval, but DCLM-7B is better by close to 10 points in LLM Foundry. In conclusion, we believe that LightEval can be potentially a good choice when evaluating smaller models, but other frameworks like LLM Foundry could give clearer signals when comparing larger models.
@@ Appendix H Hyperparameter study
A potential concern is that the training recipe can change conclusions about which dataset is optimal for training, due to interaction between training hyperparameters and dataset distributions. To address this confounder, we show in Table 13 that orderings between datasets are preserved for various combinations of weight decay and learning rate. Moreover we find that performance gains from optimal hyper-parameter choice and dataset design tend to be orthogonal and complement each other. We illustrate this effect in Table 14.
@@ Appendix I Model-based quality filters
We presented results for several different model-based quality filters in Section 4.4. In this section, we describe their implementations in further detail, focusing especially on fastText classifiers which were our method of choice for DCLM-baseline.
@@
We use the supervised fastText package from Joulin et al. [81] to train models to classify between chosen “high-quality” reference data which are given positive labels, and web-crawled data which are given negative labels. We then apply these classifiers to score each document from the pool we wish to filter, taking the predicted probability of the positive label as the score and computing a percentile-based threshold. In terms of training hyperparameters, we mostly used the default choices from the fastText package; the only hyperparameter change that we tried was to expand the feature space from unigrams only to both unigrams and bigrams (via setting the wordNgrams argument to be 2 instead of the default 1). This helped improve the quality of downstream filtered datasets, as shown in Table 15 (which extends Table 5 from Section 4.4).
@@
The bulk of our experimentation for training fastText models focused on constructing their underlying training sets, specifically the positively labeled reference data. For each experiment, we fixed the size of the training set to be 400K examples (i.e., 200K positive, 200K negative). The negatively labeled examples were sampled randomly from a set of documents that came from an earlier (smaller) version of our RefinedWeb reproduction. This version used trafilatura as the extractor instead of resiliparse, which we hypothesize might actually help for training the filtering model; as shown in Appendix J, trafilatura more aggressively removes boilerplate content that may appear in many pages (especiall from the same website). This type of content, if left in, may lead to the fastText models over-relying on these “spurious” features instead of the main contents of the page. For the positively labeled reference data, we tried several different sources, some of which involved further pre-processing:
@@
We also examined other quality filters, though found none as effective as the fastText methods described above, as shown in Table 4. We now provide further details for some of these baselines.
@@
An intuitively promising, but ultimately unfruitful approach was to consider page centrality metrics such as PageRank and Harmonic centrality metrics, with the idea that more "central" web text would yield higher quality data. We collected PageRank metrics from Common Crawl’s host level webgraph dataset [^3] https://commoncrawl.org/web-graphs and omitted any hosts that did not appear in the crawl. Next we partitioned our RefinedWeb reproduction into quintiles based on their PageRank score and trained several models at the 1B-1x scale. These results are collated in Table 16, but unfortunately no quintile performed better than a pool sampled from the union of all quintiles.
@@
         Evaluate the following paragraph for LLM pre-training suitability:      - Is it well-structured or contains useful examples to natural texts?      - Does it offer insights, useful facts or relevant information?      - Does it teach how to comply with open-ended tasks such as writing letters, poems, emails etc.?      - Is it free from harmful content?        If most criteria are met based on the content below, indicate ’yes’ for suitable. Otherwise, indicate ’no’ for unsuitable.        ### <input> ###        Is it suitable for LLM pre-training? OPTIONS:      - yes      - no
where <input> is replaced with the document tokens, clipped if too long, and appended with ‘‘…[The rest of the paragraph is omitted]’’ in such cases. While this method worked slightly better than random sampling from our pool (see Table 4), it significantly underperformed compared to our fastText experiments. Considering the high costs associated with applying it at scale, we did not perform this experiment on a larger scale.
@@
Following the success of the different deduplication methods we used (Appendix K), we studied the effect of Semantic deduplication as proposed by Abbas et al. [1]. In this approach, the authors propose embedding the documents using pre-trained language models, clustering them using k-means, and removing all but one document from each group of closely related documents to encourage diversity in the dataset. We began by embedding each document in a pool of approximately 100 million documents (following Abbas et al. [1]’s best practices) with BGE-base [176]. We then used faiss-GPU [79] to perform spherical k-means clustering, with 20 iterations and $K=11000$. We sampled documents after discarding 25% of the data. As seen in Table 4, this intervention only negatively impacted the trained model. We hypothesize that the model used for embedding has a significant impact on the outcomes of this method. However, due to the large computational overhead when scaled, making it infeasible, we opted to rely on the deduplication methods outlined in Appendix K and leave this line of research for future work.
@@
We compute basic summary statistics for each extractor based on a sample of 10 WARC files (corresponding to 900K individual pages), presenting the results in Table 17. Notably, both resiliparse and trafilatura result at least 2x shorter documents on average compared to WET files. As shown in the examples in Section J.2, WET files indeed contain many additional lines with seemingly little value for pre-training (e.g. navigation bars, boilerplate notices, copyright statements). trafilatura and resiliparse trim most of these lines out, with the former being more strict about doing so. Between the two, resiliparse still keeps in about 10% more text; some of this additional text may provide useful content such as section titles and dates for articles. In terms of runtime, the two are much farther apart, with resiliparse being roughly 8x faster.
@@
     [Trafilatura]    HERE is a sampling of some of the better antiques and flea markets around the United States.  Two or Three Times a Year  BRIMFIELD Route 20, Brimfield, Mass. 01010; 413-245-3436. Second weekend of May and July, and the second weekend after Labor Day.  RENNINGER’S OUTDOOR EXTRAVAGANZA Noble Street, Kutztown, Pa.; 717-385-0104. Thursday, Friday and Saturday of the last weekend of April, June, September.  FARMINGTON ANTIQUES WEEKEND Farmington Polo Grounds, Town Farm Road, Farmington, Conn. 06032; 508-839-9735. Starting Wednesday before shows open; 203-677-7862. June 9-10 and Sept. 1-2.  Monthly  ANN ARBOR ANTIQUES MARKET, P.O. Box 1512, Ann Arbor, Mich. 48106; 313-662-9453. May through October, third Sunday.  Continue reading the main storyKANE COUNTY FLEA MARKET, Kane County Fairgrounds, P.O. Box 549, St. Charles, Ill. 60174; 708-377-2252. Year-round, first weekend.  THE METROLINA EXPO, 7100 Statesville Road, Charlotte, N.C. 28213; 704-596-4643. Year-round, first weekend of every month.  SPRINGFIELD ANTIQUE SHOW AND FLEA MARKET, Clark County Fairgrounds, Route 41, Springfield, Ohio, 45501; 513-325-0053. Year-round, third weekend.  Weekly  BAKERSFIELD SWAP-O-RAMA, 4501 Wible Road, Bakersfield, Calif. 93313; 805-831-9342. Saturday and Sunday.  LAMBERTVILLE ANTIQUE MARKET, Route 29, Lambertville, N.J. 08530. Weekend number: 609-397-0456. Weekday: 215-752-4485, between 5 and 7 P.M. Market on Saturday and Sunday.  ATLANTA FLEA MARKET AND ANTIQUE CENTER, 5360 Peachtree Industrial Boulevard, Chamblee, Ga. 30341; 404-458-0456. Friday, Saturday and Sunday.  Continue reading the main story
@@
     [Resiliparse]    This is a digitized version of an article from The Times’s print archive, before the start of online publication in 1996. To preserve these articles as they originally appeared, The Times does not alter, edit or update them.    Occasionally the digitization process introduces transcription errors or other problems. Please send reports of such problems to archive_feedback@nytimes.com.    May 10, 1990, Page 00006 The New York Times Archives    HERE is a sampling of some of the better antiques and flea markets around the United States.    Two or Three Times a Year    BRIMFIELD Route 20, Brimfield, Mass. 01010; 413-245-3436. Second weekend of May and July, and the second weekend after Labor Day.    RENNINGER’S OUTDOOR EXTRAVAGANZA Noble Street, Kutztown, Pa.; 717-385-0104. Thursday, Friday and Saturday of the last weekend of April, June, September.    FARMINGTON ANTIQUES WEEKEND Farmington Polo Grounds, Town Farm Road, Farmington, Conn. 06032; 508-839-9735. Starting Wednesday before shows open; 203-677-7862. June 9-10 and Sept. 1-2.    Monthly    ANN ARBOR ANTIQUES MARKET, P.O. Box 1512, Ann Arbor, Mich. 48106; 313-662-9453. May through October, third Sunday.    Continue reading the main story    KANE COUNTY FLEA MARKET, Kane County Fairgrounds, P.O. Box 549, St. Charles, Ill. 60174; 708-377-2252. Year-round, first weekend.    THE METROLINA EXPO, 7100 Statesville Road, Charlotte, N.C. 28213; 704-596-4643. Year-round, first weekend of every month.    SPRINGFIELD ANTIQUE SHOW AND FLEA MARKET, Clark County Fairgrounds, Route 41, Springfield, Ohio, 45501; 513-325-0053. Year-round, third weekend.    Weekly    BAKERSFIELD SWAP-O-RAMA, 4501 Wible Road, Bakersfield, Calif. 93313; 805-831-9342. Saturday and Sunday.    LAMBERTVILLE ANTIQUE MARKET, Route 29, Lambertville, N.J. 08530. Weekend number: 609-397-0456. Weekday: 215-752-4485, between 5 and 7 P.M. Market on Saturday and Sunday.    ATLANTA FLEA MARKET AND ANTIQUE CENTER, 5360 Peachtree Industrial Boulevard, Chamblee, Ga. 30341; 404-458-0456. Friday, Saturday and Sunday.    Continue reading the main story